# SBP-FDEC: Summation-by-Parts Finite Difference Exterior Calculus

### v2

@@ 3.3 Spaces and Projections
To construct our spaces and projections, we first look at the one-dimensional case. We have a one-dimensional periodic domain $\Omega$ and a partition of $\{E^{k}\}_{k=1}^{m}$ of $\Omega$ into elements. We further assume that we have a Lagrange basis $\{\tilde{l}_{i}^{k}\}_{i=0}^{N}$ as specified in 2.1 for each element $E^{k}$, that we can get as a linear transformation of such a basis from a reference element. We further have the derived histopolation functions $\{\tilde{h}_{i}^{k}\}_{i=1}^{N}$. Now, we can construct spaces basis functions $\{l_{i}^{k}\}_{i=0,k=1}^{N-1,m}$ and $\{h_{i}^{k}\}_{i,k=1}^{N,m}$ that are defined on the whole domain as follows for all $k\in\{1,\dots,m\}$:
$l_{i}^{k}(x)$ $:=\begin{cases}\tilde{l}_{i}^{k}(x)\text{ for }x\in E^{k},\\ 0\text{ for }x\in E^{l},l\neq k,\end{cases}\qquad\forall i\in\{1,\dots,N-1\},$ (66)
$l_{0}^{k}(x)$ $:=\begin{cases}\tilde{l}_{0}^{k}(x)\text{ for }x\in E^{k},\\ \tilde{l}_{N}^{\tilde{\sigma}(k)}(x)\text{ for }x\in E^{\sigma(k)},\\ 0,\text{ for }x\in E^{l},l\neq k,l\neq{\sigma(k)}\end{cases}$ (67)
$l_{N}^{k}(x)$ $:=\begin{cases}\tilde{l}_{N}^{k}(x)\text{ for }x\in E^{k},\\ \tilde{l}_{0}^{\sigma(k)}(x)\text{ for }x\in E^{\sigma(k)},\\ 0\text{ for }x\in E^{l},l\neq k,l\neq{\sigma(k)},\end{cases}$ (68)
$h_{i}^{k}(x)$ $:=\begin{cases}\tilde{h}_{i}^{k}(x)\text{ for }x\in E^{k},\\ 0\text{ for }x\in E^{l},l\neq k,\end{cases}\qquad\forall i\in\{1,\dots,N\},$ (69)
where $\sigma(k)$ is defined as the element index of the right neighbor of the element with index $k$ and $\tilde{\sigma}(k)$ is the index of the left neighbor of element $k$. Here $l_{N}^{k}$ and $l_{0}^{\sigma(k)}$ refer to the same function, as do $l_{0}^{k}$ and $l_{N}^{\tilde{\sigma}(k)}$. While the Lagrange basis is globally continuous, the histopolation basis does not have a unique value at element boundaries. When we use the nodal evaluation matrix for the histopolation functions (15), we will do so for each element separately and obtain the value of the continuous extension of the basis functions to the boundary when restricted to the element in question.
Analogously to 2.1 we can now define the space $U_{l}$ as the span of the global Lagrange functions and $U_{h}$ as the span of the global histopolation functions. We again have $\frac{\partial{}}{\partial{x}}(U_{l})=U_{h}$, where $\frac{\partial{}}{\partial{x}}$ is the weak derivative operator, even though we have in the periodic case the same number of degrees of freedom. This is because there is an additional continuity requirement at the domain boundary that breaks down when taking the derivative. Now, in analogy to (8) and (9), let $p^{0}$ be the global histopolation operator and $p^{1}$ the global interpolation operator associates with the defined bases.
We can apply the derivative formula (12) element-wise to obtain the commutation property (10) for the multi-element case. This means the commuting diagram 1 for the one-dimensional de Rham complex commutes. To extend this approach to 2D, we use a tensor product ansatz, where for a rectangular periodic domain $\Omega$ with a Cartesian grid, we define the discrete spaces via a tensor product ansatz as follows:
$W^{0}$ $:=U_{h}\otimes U_{h},$ (70)
@@
$W^{2}$ $:=U_{l}\otimes U_{l}.$ (72)
Here we have a slight abuse of notation because we can have different number of elements and scaling factors in both directions. The projections $p_{0},\mathbf{p}_{1},$ and $p_{2}$ are now defined by taking the interpolation in the directions where we use the space $U_{l}$ and histopolation in the directions where we have $U_{h}$:
$p^{0}(f)$ $:=\sum_{k=1}^{m}\sum_{i,j=1}^{N}\sum_{k=1}^{m}\sum_{i,j=1}^{N}\left(\int_{x_{i-1}^{k}}^{x_{i}^{k}}\int_{y_{j-1}^{k}}^{y_{j}^{k}}f(s,t)\text{ d}s\text{ d}t\right)h_{i}^{k}(x)h_{j}^{k}(y),$ (73)
@@
$p^{2}(f)$ $:=\sum_{k=1}^{m}\sum_{i,j=0}^{N-1}f(x_{i}^{k},y_{j}^{k})l_{i}(x)l_{j}(y).$ (76)
Since we use the same DOFs as Gerritsma et al. [gerritsma_preprint], the boundedness of the projections carries over, since we can factorize over Gerritsma’s projections switching out the basis functions. As a linear mapping between finite dimensional vector spaces, it is bounded, and therefore our projections are bounded as the composition of two bounded mappings. It is straightforward to see that we also get commuting property for the discrete 2D de Rham complex 5.
@@ 4.1 Galerkin Discretization
$\underline{B}^{z}=\begin{pmatrix}B^{z}_{0,0;1}\\ \vdots\\ B^{z}_{N-1,0;1}\\ B^{z}_{0,1;1}\\ \vdots\\ B^{z}_{N-1,N-1;1}\\ B^{z}_{0,0;2}\\ \vdots\\ B^{z}_{N-1,N-1;m}\end{pmatrix},\qquad\underline{E}^{x}=\begin{pmatrix}\overline{E}^{x}_{0,1;1}\\ \vdots\\ \overline{E}^{x}_{N-1,1;1}\\ \overline{E}^{x}_{0,2;1}\\ \vdots\\ \overline{E}^{x}_{N-1,N;1}\\ \overline{E}^{x}_{0,1;2}\\ \vdots\\ \overline{E}^{x}_{N-1,N;m}\end{pmatrix},\qquad\underline{E}^{x}=\begin{pmatrix}\overline{E}^{y}_{1,0;1}\\ \vdots\\ \overline{E}^{y}_{N,0;1}\\ \overline{E}^{y}_{1,1;1}\\ \vdots\\ \overline{E}^{y}_{N,N-1;1}\\ \overline{E}^{y}_{1,0;2}\\ \vdots\\ \overline{E}^{y}_{N,N-1;m}\end{pmatrix}$.
For ease of notation, we define $B_{N,j,k}:=B_{0,j,\sigma_{r}(k)}$, $B_{i,N,k}:=B_{i,0,\sigma_{r}(k)}$, and $B_{N,N,k}:=B_{0,0,\sigma_{r}\left(\sigma_{u}(k)\right)}$. Here, $\sigma_{r}(k)$ is defined as the index of the right neighbor element of the element $E^{k}$ and $\sigma_{u}(k)$ is the index of the upper neighbor of $E^{k}$. We also apply this notation for the electric field coefficients in directions where the electric field component is continuous. Additionally, we assume for simplicity that the reference grid of our finite difference operator is of length one (which we can achieve by rescaling). Further, we assume that each element has the same length $\Delta x$ in $x$- and $\Delta y$ in $y$-direction. Both assumptions are not necessary, they are used to simplify the definition of our matrix operators. Since we have Lagrange functions associated with boundary nodes whose support contains two elements, we define intertwined quadrature weights
@@
We can now derive our method via the Galerkin ansatz in (63) and (64). Using our basis functions as test functions we get for the first component of (64):
$\begin{split}&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=1}^{N}\int_{E^{k}}\frac{\partial{}}{\partial{t}}(E^{x}_{i,j;k})l^{k}_{i}(x)h^{k}_{j}(x)l^{l}_{s}(x)h^{l}_{t}(y)\text{ d}x\text{ d}y\\ =&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=0}^{N-1}\int_{E^{k}}\frac{\partial{}}{\partial{y}}\left(B^{z}_{i,j;k}l^{k}_{i}(x)l^{k}_{j}(y)\right)l^{l}_{s}(x)h^{l}_{t}(y)\text{ d}x\text{ d}y\\ =&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=1}^{N}\int_{E^{k}}\left(B^{z}_{i,j;k}-B^{z}_{i,j-1;k}\right)l^{k}_{i}(x)h^{k}_{j}(y)l^{l}_{s}(x)h^{l}_{t}(y)\text{ d}x\text{ d}y.\end{split}$ (78)
@@
$\begin{split}&\Delta x\Delta y\tilde{\omega}_{s}\sum_{j=1}^{N}\frac{\partial{}}{\partial{t}}(E^{x}_{s,j;l})\sum_{o=0}^{N}\omega_{o}h_{j}^{k}(y^{l}_{o})h^{l}_{t}(y^{l}_{o})\\ =&\Delta x\Delta y\tilde{\omega}_{s}\sum_{j=1}^{N}\left(B^{z}_{s,j;l}-B^{z}_{s,j-1;l}\right)\sum_{o=0}^{N}\omega_{o}h_{j}^{k}(y^{l}_{o})h^{l}_{t}(y^{l}_{o}),\end{split}$ (79)
@@
$\underline{V}^{T}\underline{M}\underline{V}\underline{E}_{s;l}^{x}=\underline{V}^{T}\underline{M}\underline{V}\underline{\Delta}\underline{B}_{s;l}^{z},$ (80)
with $\underline{E}_{s;l}^{x}:=\left(\frac{\partial{}}{\partial{t}}(E^{x}_{s,1;l}),\dots\frac{\partial{}}{\partial{t}}(E^{x}_{s,N;l})\right)^{T}$ and $\underline{B}_{s;l}^{z}:=\left(B^{x}_{s,0;l},\dots E^{x}_{s,N;l}\right)^{T}$. Usually the matrix $\underline{V}$ would contain a scaling factor $(\Delta x)^{-1}$ from the derivative matrix, but this cancels out as well. Since the histopolation mass matrix is invertible, we obtain for all element indices $l\in\{1,\dots,m\}$
$\frac{\partial{}}{\partial{t}}(E^{x}_{s,j;l})=B^{z}_{s,j;l}-B^{z}_{s,j-1;l},\qquad\forall s\in\{0,\dots,N\},j\in\{1,\dots,N\}.$ (81)
@@
$\frac{\partial{}}{\partial{t}}(E^{y}_{i,t;l})=B^{z}_{i,t;l}-B^{z}_{i-1,t;l},\qquad\forall t\in\{0,\dots,N\},i\in\{1,\dots,N\}.$ (82)
@@
$\begin{split}&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=0}^{N-1}\int_{E^{k}}\frac{\partial{}}{\partial{t}}(B^{x}_{i,j;k})l^{k}_{i}(x)l^{k}_{j}(x)l^{l}_{s}(x)l^{l}_{t}(y)\text{ d}x\text{ d}y\\ =&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=1}^{N}\int_{E^{k}}E^{x}_{i,j;k}l^{k}_{i}(x)h^{k}_{j}(y)l^{l}_{s}(x)\frac{\partial{}}{\partial{y}}l^{l}_{t}(y)\text{ d}x\text{ d}y\\ -&\sum_{k=1}^{m}\sum_{i=1}^{N}\sum_{j=0}^{N-1}\int_{E^{k}}E^{y}_{i,j;k}h^{k}_{i}(x)l^{k}_{j}(y)\frac{\partial{}}{\partial{x}}l^{l}_{s}(x)l^{l}_{t}(y)\text{ d}x\text{ d}y.\end{split}$ (83)
@@
$\begin{split}&\tilde{\omega}_{s}\tilde{\omega}_{t}\frac{\partial{}}{\partial{t}}(B^{x}_{s,t;l})\\ =&\tilde{\omega}_{s}\sum_{j=1}^{N}E^{x}_{s,j;l}\sum_{o=0}^{N}\omega_{o}h^{l}_{j}(y_{o}^{l})\frac{\partial{}}{\partial{y}}l^{l}_{t}(y_{o}^{l})+T_{t;l}\\ -&\tilde{\omega}_{t}\sum_{i=1}^{N}E^{y}_{i,t;l}\sum_{o=0}^{N}\omega_{o}h^{l}k_{i}(x_{o}^{l})\frac{\partial{}}{\partial{x}}l^{l}_{s}(x_{o}^{l})+S_{s;l},\end{split}\qquad\forall s,t\in\{0,\dots,N\}.$ (84)
@@
$T_{t;l}=\tilde{\omega}_{s}\sum_{j=1}^{N}E^{x}_{s,j;\sigma_{u}(l)}\sum_{o=0}^{N}\omega_{o}h^{\sigma_{u}(l)}_{j}(y_{o}^{\sigma_{u}(l)})\frac{\partial{}}{\partial{y}}l^{l}_{t}(y_{o}^{\sigma_{u}(l)}).$ (85)
@@ 4.2 Matrix-Vector Formulation
$\underline{\hat{\Delta}}$ $:=\left(\delta_{i+1,j}-\delta_{i,j}\right)_{i=0,j=0}^{N-1},$ (90)
$\underline{\tilde{\Delta}}$ $:=\left(\delta_{i,N-1}\delta_{j,0}\right)_{i=0,j=0}^{N-1}.$ (91)
Furthermore, we have the Vandermonde matrix $\underline{V}\in\mathbb{R}^{N+1\times N}$ defined in (15). To construct the $2D$ operators we need the neighbour matrices $\underline{P}^{r}\in\mathbb{R}^{m\times m}$ and $\underline{P}^{u}\in\mathbb{R}^{m\times m}$ that encode the neighbour information as follows:
@@
$\underline{\Delta}^{2D}_{y}$ $:=\underline{I}_{m}\otimes\hat{\underline{\Delta}}\otimes\underline{I}_{N}+\underline{P}_{u}\otimes\tilde{\underline{\Delta}}\otimes\underline{I}_{N},$ (101)
@@
$\underline{V}_{x}$ $:=\left(\Delta x\right)^{-1}\left(\underline{I}_{m}\otimes\underline{I}_{N}\otimes\underline{V}\right),$ (103)
$\underline{V}_{y}$ $:=\left(\Delta y\right)^{-1}\left(\underline{I}_{m}\otimes\underline{V}\otimes\underline{I}_{N}\right).$ (104)
@@
Using the Summation-by-Parts property we can also derive a strong form version of the scheme. First we rewrite the semidiscrete Faraday equation (109), using the identity $\underline{V}\underline{\Delta}=\underline{D}$
@@
We note that on a function level, the derivative matrix $\underline{D}$ computes the nodal derivatives of the element-local Lagrange functions, not the histopolation functions. Since we take the derivative in the directions we have the histopolation basis functions, this amounts to interpolating the functions with the element-local Lagrange basis, so without continuity requirement, and then computing the local derivative.
The strong form equation (112) then consists of the element-local interpolation derivative in x-direction for $E^{y}$ and in y-direction for $E^{x}$ for $k\neq 0$ and $l\neq 0$ respectively. If the degree of freedom in question is on the element boundary in x- or y-direction, we instead have a weighted jump penalty term minus the weighted average of the discontinuous derivative at the boundary. This results in a treatment of the discontinuous derivative similar to a finite difference SAT, a finite volume or a discontinuous Galerkin ansatz.
@@
since $\underline{J}$ is anti-symmetric.

### v3

@@ 3.3 Spaces and Projections
To construct our spaces and projections, we first look at the one-dimensional case. We have an interval $[a,b]$ and a partition of $\{K^{k}\}_{k=1}^{m}$ of a,b{a,b} into elements. The global $U_{l}$ and $U_{h}$ spaces are defined via the local $U_{l}$ and $U_{h}$ spaces on each element. In the case of $U_{l}$ we impose global continuity to guarantee $W^{0}\subset H^{1}$ and $W^{1}\subset H(\text{div})$ (see for example Monk [32]), while there are no such requirements for $U_{h}$. Further details can be found in appendix C. To extend this approach to 2D, we use a tensor product ansatz, where for a rectangular periodic domain $\Omega$ with a Cartesian grid, we define the discrete spaces via a tensor product ansatz as follows:
$W^{0}$ $:=U_{l}\otimes U_{l},$ (83)
@@
$W^{2}$ $:=U_{h}\otimes U_{h}.$ (85)
Here we have a slight abuse of notation because we can have different number of elements and scaling factors in both directions. The projections $p^{0},\mathbf{p}^{1},$ and $p^{2}$ are now defined by taking the interpolation in the directions where we use the space $U_{l}$ and histopolation in the directions where we have $U_{h}$:
$p^{0}(f)$ $:=\sum_{k=1}^{m}\sum_{i,j=0}^{N-1}f(x_{i}^{k},y_{j}^{k})l_{i}(x)l_{j}(y),$ (86)
@@
$p^{2}(f)$ $:=\sum_{k=1}^{m}\sum_{i,j=1}^{N}\left(\int_{x_{i-1}^{k}}^{x_{i}^{k}}\int_{y_{j-1}^{k}}^{y_{j}^{k}}f(s,t)\text{ d}s\text{ d}t\right)h_{i}^{k}(x)h_{j}^{k}(y),$ (89)
where $m$ is the number of elements. Since we use the same DOFs as Gerritsma et al. [27], the boundedness of the projections carries over, since we can factorize over Gerritsma’s projections switching out the basis functions. As a linear mapping between finite dimensional vector spaces, it is bounded, and therefore our projections are bounded as the composition of two bounded mappings. It is straightforward to see that we also get commuting property for the discrete 2D de Rham complex in Fig. 5.
@@ 4.1 Galerkin Discretization
$\underline{B}^{z}=\begin{pmatrix}B^{z}_{0,0;1}\\ \vdots\\ B^{z}_{N-1,0;1}\\ B^{z}_{0,1;1}\\ \vdots\\ B^{z}_{N-1,N-1;1}\\ B^{z}_{0,0;2}\\ \vdots\\ B^{z}_{N-1,N-1;m}\end{pmatrix},\qquad\overline{\underline{E}}^{x}=\begin{pmatrix}\overline{E}^{x}_{0,1;1}\\ \vdots\\ \overline{E}^{x}_{N-1,1;1}\\ \overline{E}^{x}_{0,2;1}\\ \vdots\\ \overline{E}^{x}_{N-1,N;1}\\ \overline{E}^{x}_{0,1;2}\\ \vdots\\ \overline{E}^{x}_{N-1,N;m}\end{pmatrix},\qquad\overline{\underline{E}}^{x}=\begin{pmatrix}\overline{E}^{y}_{1,0;1}\\ \vdots\\ \overline{E}^{y}_{N,0;1}\\ \overline{E}^{y}_{1,1;1}\\ \vdots\\ \overline{E}^{y}_{N,N-1;1}\\ \overline{E}^{y}_{1,0;2}\\ \vdots\\ \overline{E}^{y}_{N,N-1;m}\end{pmatrix}$.
For ease of notation, we define $B_{N,j,k}:=B_{0,j,\sigma_{r}(k)}$, $B_{i,N,k}:=B_{i,0,\sigma_{r}(k)}$, and $B_{N,N,k}:=B_{0,0,\sigma_{r}\left(\sigma_{u}(k)\right)}$. In the non-periodic case, the vector segments for the right and upper boundary elements contain the boundary value as well. Here, $\sigma_{r}(k)$ is defined as the index of the right neighbor element of the element $K_{k}$ and $\sigma_{u}(k)$ is the index of the upper neighbor of $K_{k}$. We also apply this notation for the electric field coefficients in directions where the electric field component is continuous. Additionally, we assume for simplicity that the reference grid of our finite difference operator is of length one (which we can achieve by rescaling). Further, we assume that each element has the same length $\Delta x$ in $x$- and $\Delta y$ in $y$-direction. Both assumptions are not necessary, they are used to simplify the definition of our matrix operators. Since we have Lagrange functions associated with boundary nodes whose support contains two elements, we define intertwined quadrature weights
@@
We can now derive our method via the Galerkin ansatz in (80) and (81). In this section we derive the scheme for the periodic case. The non-periodic case is almost the same, but one has to adjust the index range of the sums for boundary elements and account for the fact, that some elements have no neighbour in one or more directions. Using our basis functions as test functions we get for the first component of (81):
$\begin{split}&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=1}^{N}\int_{K_{k}}\frac{\partial{}}{\partial{t}}(\overline{E}^{x}_{i,j;k})l^{k}_{i}(x)h^{k}_{j}(x)l^{l}_{s}(x)h^{l}_{t}(y)\text{ d}x\text{ d}y\\ =&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=0}^{N-1}\int_{K_{k}}\frac{\partial{}}{\partial{y}}\left(B^{z}_{i,j;k}l^{k}_{i}(x)l^{k}_{j}(y)\right)l^{l}_{s}(x)h^{l}_{t}(y)\text{ d}x\text{ d}y\\ =&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=1}^{N}\int_{K_{k}}\left(B^{z}_{i,j;k}-B^{z}_{i,j-1;k}\right)l^{k}_{i}(x)h^{k}_{j}(y)l^{l}_{s}(x)h^{l}_{t}(y)\text{ d}x\text{ d}y.\end{split}$ (91)
@@
$\begin{split}&\Delta x\Delta y\tilde{\omega}_{s}\sum_{j=1}^{N}\frac{\partial{}}{\partial{t}}(\overline{E}^{x}_{s,j;l})\sum_{o=0}^{N}\omega_{o}h_{j}^{k}(y^{l}_{o})h^{l}_{t}(y^{l}_{o})\\ =&\Delta x\Delta y\tilde{\omega}_{s}\sum_{j=1}^{N}\left(B^{z}_{s,j;l}-B^{z}_{s,j-1;l}\right)\sum_{o=0}^{N}\omega_{o}h_{j}^{k}(y^{l}_{o})h^{l}_{t}(y^{l}_{o}),\end{split}$ (92)
@@
$\underline{V}^{T}\underline{M}\underline{V}\frac{d}{dt}\underline{\overline{E}}_{s;l}^{x}=\underline{V}^{T}\underline{M}\underline{V}\underline{\Delta}\underline{B}_{s;l}^{z},$ (93)
with $\frac{d}{dt}\underline{\overline{E}}_{s;l}^{x}:=\left(\frac{d}{dt}(\overline{E}^{x}_{s,1;l}),\dots\frac{d}{dt}(\overline{E}^{x}_{s,N;l})\right)^{T}$ and $\underline{B}_{s;l}^{z}:=\left(B^{z}_{s,0;l},\dots B^{z}_{s,N;l}\right)^{T}$. Usually the matrix $\underline{V}$ would contain a scaling factor $(\Delta x)^{-1}$ from the derivative matrix, but this cancels out as well. Since the histopolation mass matrix is invertible, we obtain for all element indices $l\in\{1,\dots,m\}$
$\frac{\partial{}}{\partial{t}}(\overline{E}^{x}_{s,j;l})=B^{z}_{s,j;l}-B^{z}_{s,j-1;l},\qquad\forall s\in\{0,\dots,N\},j\in\{1,\dots,N\}.$ (94)
@@
$\frac{\partial{}}{\partial{t}}(\overline{E}^{y}_{i,t;l})=B^{z}_{i,t;l}-B^{z}_{i-1,t;l},\qquad\forall t\in\{0,\dots,N\},i\in\{1,\dots,N\}.$ (95)
@@
$\begin{split}&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=0}^{N-1}\int_{K_{k}}\frac{\partial{}}{\partial{t}}(B^{x}_{i,j;k})l^{k}_{i}(x)l^{k}_{j}(x)l^{l}_{s}(x)l^{l}_{t}(y)\text{ d}x\text{ d}y\\ =&\sum_{k=1}^{m}\sum_{i=0}^{N-1}\sum_{j=1}^{N}\int_{K_{k}}\overline{E}^{x}_{i,j;k}l^{k}_{i}(x)h^{k}_{j}(y)l^{l}_{s}(x)\frac{\partial{}}{\partial{y}}l^{l}_{t}(y)\text{ d}x\text{ d}y\\ -&\sum_{k=1}^{m}\sum_{i=1}^{N}\sum_{j=0}^{N-1}\int_{K_{k}}\overline{E}^{y}_{i,j;k}h^{k}_{i}(x)l^{k}_{j}(y)\frac{\partial{}}{\partial{x}}l^{l}_{s}(x)l^{l}_{t}(y)\text{ d}x\text{ d}y.\end{split}$ (96)
@@
$\begin{split}&\tilde{\omega}_{s}\tilde{\omega}_{t}\frac{\partial{}}{\partial{t}}(B^{x}_{s,t;l})\\ =&\tilde{\omega}_{s}\sum_{j=1}^{N}\overline{E}^{x}_{s,j;l}\sum_{o=0}^{N}\omega_{o}h^{l}_{j}(y_{o}^{l})\frac{\partial{}}{\partial{y}}l^{l}_{t}(y_{o}^{l})+T_{t;l}\\ -&\tilde{\omega}_{t}\sum_{i=1}^{N}\overline{E}^{y}_{i,t;l}\sum_{o=0}^{N}\omega_{o}h^{l}_{i}(x_{o}^{l})\frac{\partial{}}{\partial{x}}l^{l}_{s}(x_{o}^{l})+S_{s;l},\end{split}\qquad\forall s,t\in\{0,\dots,N\}.$ (97)
@@
$T_{t;l}=\tilde{\omega}_{s}\sum_{j=1}^{N}\overline{E}^{x}_{s,j;\sigma_{u}(l)}\sum_{o=0}^{N}\omega_{o}h^{\sigma_{u}(l)}_{j}(y_{o}^{\sigma_{u}(l)})\frac{\partial{}}{\partial{y}}l^{l}_{t}(y_{o}^{\sigma_{u}(l)}).$ (98)
@@ 4.2 Matrix-Vector Formulation
$\underline{\hat{\Delta}}$ $:=\left(\delta_{i+1,j}-\delta_{i,j}\right)_{i=0,j=0}^{N-1,N-1},$ (103)
$\underline{\tilde{\Delta}}$ $:=\left(\delta_{i,N-1}\delta_{j,0}\right)_{i=0,j=0}^{N-1,N-1}.$ (104)
Here $(a_{i,j})_{i=0,j=0}^{K,L}$ is defined as the $\mathbb{R}^{K\times L}$ matrix with entries $a_{i,j}$ with row indices running from $i=0$ to $K$ and column indices $j=0$ to $L$. Since we assume $\frac{\partial{}}{\partial{x}}l_{j}(x_{i})=D_{i,j}$ for the derivative matrix of the SBP operator from (62) these split derivative matrices just take their value from the derivative matrix of the SBP operator. Furthermore, we have the Vandermonde matrix $\underline{V}\in\mathbb{R}^{N+1\times N}$ defined in (15). To construct the $2D$ operators we need the neighbour matrices $\underline{P}^{r}\in\mathbb{R}^{m\times m}$ and $\underline{P}^{u}\in\mathbb{R}^{m\times m}$ that encode the neighbour information as follows:
@@
$\underline{\Delta}^{2D}_{y}$ $:=\underline{I}_{m}\otimes\hat{\underline{\Delta}}\otimes\underline{I}_{N}+\underline{P}_{u}\otimes\tilde{\underline{\Delta}}\otimes\underline{I}_{N}.$ (114)
@@
$\underline{V}^{2D}_{x}$ $:=\left(\Delta x\right)^{-1}\left(\underline{I}_{m}\otimes\underline{I}_{N}\otimes\underline{V}\right),$ (116)
$\underline{V}^{2D}_{y}$ $:=\left(\Delta y\right)^{-1}\left(\underline{I}_{m}\otimes\underline{V}\otimes\underline{I}_{N}\right).$ (117)
@@
If we want to impose homogeneous boundary conditions, we have to change the matrix operators slightly. We need versions for the intertwined mass matrix (100) and the difference matrix (17) for the left/lower and the right/upper boundary elements. The intertwined mass matrix for the boundary elements read
$\underline{\hat{M}}_{l}$ $=\text{diag}\left(\omega_{0},\dots,\omega_{N-1}\right),$ (123)
$\underline{\hat{M}}_{r}$ $=\text{diag}\left(\omega_{0}+\omega_{N},\dots,\omega_{N-1},\omega_{N}\right),$ (124)
$\underline{\hat{M}}_{1}$ $=\text{diag}\left(\omega_{0},\dots,\omega_{N-1},\omega_{N}\right),$ (125)
where $\underline{\hat{M}}_{1}$ is the intertwined mass matrix in the case of one element. The boundary replacements for the cutoff difference matrices $\hat{\underline{\Delta}}$ read
$\underline{\hat{\Delta}}_{l}$ $=\underline{\hat{\Delta}}$, (126)
$\underline{\hat{\Delta}}_{r}$ $=\underline{\Delta},$ (127)
$\underline{\hat{\Delta}}_{1}$ $=\underline{\Delta}.$ (128)
The remaining part $\underline{\tilde{\Delta}}$ of the difference matrix entries remain the same, but the dimensions have to be adjusted to match the number of DOFs in the boundary elements by adding additional zeros.
The 2D matrices do no longer have the form of a Kronecker product over all elements. The mass and Vandermonde matrices are still block-diagonal matrices with one block for each element, but since we have to adjust for the boundary elements, we cannot take a Kronecker product with $I_{m}$ anymore. For the 2D Vandermonde matrix the only thing that changes for different blocks is the size of the identity matrix in the Kronecker product. For example, if we have an element $K$ on the upper boundary, then the block for $V_{x}$ for this element reads
$V^{2D}_{x,K}=\underline{I}_{N+1}\otimes\underline{V}.$ (129)
For the 2D mass and Vandermonde matrices we have to use the boundary versions for $\hat{M}$, $\underline{\hat{\Delta}}$ and $\underline{\tilde{\Delta}}$. The mass matrix $\underline{M}^{2D}$, that only uses element-local mass-matrices, remains unchanged. If we want to impose essential boundary conditions, we have to truncate the boundary matrices, since the test functions corresponding to boundary DOFs are no longer part of the test space. Then we have
$\underline{\hat{M}}_{l}$ $=\text{diag}\left(\omega_{1},\dots,\omega_{N-1}\right),$ (130)
$\underline{\hat{M}}_{r}$ $=\text{diag}\left(\omega_{0}+\omega_{N},\dots,\omega_{N-1}\right),$ (131)
$\underline{\hat{M}}_{1}$ $=\text{diag}\left(\omega_{1},\dots,\omega_{N-1},\omega_{N}\right),$ (132)
and
$\underline{\hat{\Delta}}_{l}$ $=\underline{\hat{\Delta}}_{1:N,2:N+1},$ (133)
$\underline{\hat{\Delta}}_{r}$ $=\underline{\Delta}_{1:N,1:N},$ (134)
$\underline{\hat{\Delta}}_{1}$ $=\underline{\Delta}_{1:N,2:N},$ (135)
where $\underline{\Delta}_{1:N,1:N}$ is the matrix we get when we take all rows, but only the first $N$ columns from $\underline{\Delta}$. Alternatively one can use the untruncated versions and set the boundary DOFs to 0 after each right-hand side evaluation, which is the variant that is implemented in the linked github repository. Using the Summation-by-Parts property we can also derive a strong form version of the scheme. First we rewrite the semi-discrete Faraday equation (122), using the identity $\underline{V}\underline{\Delta}=\underline{D}$
@@
We note that on a function level, the derivative matrix $\underline{D}$ computes the nodal derivatives of the element-local Lagrange functions, not the histopolation functions. Since we take the derivative in the directions we have the histopolation basis functions, this amounts to interpolating the functions with the element-local Lagrange basis, without requiring continuity, and then computing the local derivative. The strong form equation (138) then consists of the element-local interpolation derivative in x-direction for $E^{y}$ and in y-direction for $E^{x}$ for $k\neq 0$ and $l\neq 0$ respectively. If the degree of freedom in question is on the element boundary in x- or y-direction, we instead have a weighted jump penalty term minus the weighted average of the discontinuous derivative at the boundary. This results in a treatment of the discontinuous derivative similar to a finite difference SAT, a finite volume or a discontinuous Galerkin ansatz.
@@
since $\underline{J}$ is anti-symmetric. We remark that this property can only be proven for the weak form of the scheme. For non-SBP operators strong and weak form are not equivalent, so this may not hold for the strong form of the scheme for non-SBP operators.